OpenStudy (anonymous):

The base of a triangle is 5 cm greater than the height. The area is 52 cm^2. Find the height and the length of the base.

hero (hero):

52 = 1/2(x)(x+5)

hero (hero):

104 = (x)(x+5)

hero (hero):

0 = x^2 + 5x - 104

OpenStudy (anonymous):

I am trying to figure this out but not getting very far yet.

hero (hero):

x = 8 x+5 = 13

hero (hero):

I skipped a whole bunch of steps, but that's your answer

OpenStudy (anonymous):

LOL, I got -13 and 8 and I knew that couldn't be right. So the base is 13 and the height is 8?

hero (hero):

Yes

OpenStudy (anonymous):

great! thank you.

hero (hero):

yeah, -13 is negative, so it wouldn't work. Which means only one x would work. You add 5 to that and you get positive 13

OpenStudy (anonymous):

thanks!